A carton of whipping cream contains 35% butterfat. A carton of "2%" milk contains 2% butterfat. If a volume of whipping cream is mixed with an equal volume of "2%" milk, the percentage of butterfat in the mixture?
Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(75887)

A carton of whipping cream contains 35% butterfat. A carton of "2%" milk contains 2% butterfat. If a volume of whipping cream is mixed with an equal volume of "2%" milk, the percentage of butterfat in the mixture?
-----
fat + fat = fat
0.35x + 0.02x = p*(2x)
0.37x = 2px
0.37 = 2p
percentage = 0.185 = 18.5%
==============================
Cheers,
Stan H.
=============

Answer by Alan3354(69443)

A carton of whipping cream contains 35% butterfat. A carton of "2%" milk contains 2% butterfat. If a volume of whipping cream is mixed with an equal volume of "2%" milk, the percentage of butterfat in the mixture?
------------
It will be the average
(2 + 35)/2 = 18.5%